# ParamISP: Learned Forward and Inverse ISPs using Camera Parameters

### v1

@@ Abstract
RAW images are rarely shared mainly due to its excessive data size compared to their sRGB counterparts obtained by camera ISPs. Learning the forward and inverse processes of camera ISPs has been recently demonstrated, enabling physically-meaningful RAW-level image processing on input sRGB images. However, existing learning-based ISP methods fail to handle the large variations in the ISP processes with respect to camera parameters such as ISO and exposure time, and have limitations when used for various applications. In this paper, we propose ParamISP, a learning-based method for forward and inverse conversion between sRGB and RAW images, that adopts a novel neural-network module to utilize camera parameters, which is dubbed as ParamNet. Given the camera parameters provided in the EXIF data, ParamNet converts them into a feature vector to control the ISP networks. Extensive experiments demonstrate that ParamISP achieve superior RAW and sRGB reconstruction results compared to previous methods and it can be effectively used for a variety of applications such as deblurring dataset synthesis, raw deblurring, HDR reconstruction, and camera-to-camera transfer.
@@ 1 Introduction
Unlike sRGB images, RAW images provide physically-meaningful and interpretable information such as noise distributions as they have the linear relationship between image intensity and radiant energy incident on a camera sensor. Such properties of RAW images have been exploited for denoising [1, 17, 23], HDR [5, 14], and super-resolution [32, 24], leading to superior reconstruction quality than using only sRGB images. However, a RAW image demands large memory to store due to the use of per-pixel high-precision bits without any compression. Therefore, only sRGB images are often shared without their RAW counterparts, making it difficult to utilize the useful properties of RAW images.
Recent methods have studied ways of reconstructing RAW images from sRGB images and vice versa by modeling forward and inverse ISPs [4, 30, 29, 3, 7]. However, despite notable improvements, existing methods suffer from limitations. First, they do not account for the adaptive nature of the real-world ISPs, which adjust their operations based on the camera parameters such as exposure time, sensor sensitivity and focal length accessible in the EXIF metadata of a JPEG image. The operations of real-world ISPs such as denoising, sharpening and tone adjustment depend on the camera parameters, producing images with different tones and colors, as shown in Fig. 1. Furthermore, previous methods adopt rather simple network architectures disregarding the complexity of the ISP operations, which leads to low RAW reconstruction quality. They construct a single network by stacking residual blocks [30] or invertible blocks [29], or organize the modules by simply arranging convolution layers [3] continuously.
@@
In this paper, we present a novel forward and inverse ISP framework, ParamISP. ParamISP is designed to faithfully reflect the real-world ISP operations that change based on camera parameters. To effectively incorporate the camera parameters, ParamISP consists of a pair of forward and inverse ISP networks that include a novel neural network module ParamNet. ParamNet extracts a feature vector from the camera parameters including exposure time, sensitivity, aperture size, and focal length, and feeds it to the forward and inverse ISP subnetworks to control their behaviors.
@@
    We propose ParamISP, a novel learning-based forward and inverse ISP framework that leverages camera parameters (e.g., exposure time, sensitivity, aperture size, and focal length) for high-quality sRGB and RAW reconstruction.
@@ 3.2 Forward ISP Network
GlobalNet globally adjusts tone and color according to the content of its input image and the optical parameter feature vector $z$. To this end, GlobalNet adopts parametric global adjustment operations and predicts their parameters. Specifically, GlobalNet adopts two types of global adjustment operations: quadratic transformation $f_{q}$ and gamma correction $f_{g}$. Let us denote the input image to global adjustment operation as $I$, and its pixel as $p=(p_{r},p_{g},p_{b})$. Then, the quadratic transformation $f_{q}$ is defined as:
@@
where $f_{gq}^{n}(I)$ is defined as $f_{gq}^{n}=f_{g}^{n}(f_{q}^{n}(I))$. $f_{g}^{n}$ and $f_{q}^{n}$ are the gamma correction and quadratic transformation with the $n$-th coefficients $G^{n}$ and $W^{n}$, respectively. In our experiments, we use $N=4$ as it leads to the best quality (Tab. 4).
@@ 4.1 Ablation Study
ISP Network Architecture
Additionally, we verify the effects of a series of gamma correction and quadratic transformation in GlobalNet (Tab. 4). We compare global adjustment operations of previous methods [3, 23] to GlobalNet with a series of gamma correction and quadratic transformation (1st to 3rd columns). Subsequently, we progressively increase the number of gamma correction and quadratic transformation in the series to validate the effects of overlaying (4th to 7th columns). The table shows that our global tone adjustment operation improves the PSNR by more than 0.89 dB over the others [3, 23].
@@
  Operation $A$ $B$ $C$$\times$1 $C$$\times$2 $C$$\times$3 $C$$\times$4 $C$$\times$5
[…]
Table 4: Ablation study on the effects of global adjustment operation. $A$ and $B$ represent 3$\times$6 polynomial mapping function (CIE XYZ Net [3]) and 3$\times$10 quadratic transformation (DeepISP [23]) respectively, while $C$ denotes a series of gamma correction and quadratic transformation (Ours). Additionally, in $C$$\times$$N$, $N$ represents the number of pairs ($N=1,2,3,4,5$).
@@ 5 Applications
Deblurring in RAW Space
Figure 8: Qualitative results of HDR reconstruction.
@@
Fig. 8 shows qualitative results using different models. Previous methods primarily focus on cyclic reconstruction. Specifically, CycleISP uses an input LDR sRGB image for the forward ISP to adjust the tone, while InvISP, composed of a single flow-based network, is not robust when intermediate RAW images are altered. On the other hand, ParamISP successfully produces an HDR reconstruction result compared to other methods.
@@ 6 Conclusion
In this paper, we present ParamISP, a novel learning-based forward and inverse ISP framework that leverages camera parameters (e.g., exposure time, sensitivity, aperture size, and focal length). To incorporate camera parameters effectively, we introduce ParamNet to control the forward and inverse ISP networks, proposing a stable and effective training strategy with respect to camera parameters. We also present novel network architectures for ISP networks that better reflect real-world ISP operations. Through our extensive experiments on different cameras, ParamISP shows the state-of-the-art performance in RAW and sRGB reconstruction and the robust applicability to various applications compared to previous methods.

### v2

@@ Abstract
RAW images are rarely shared mainly due to its excessive data size compared to their sRGB counterparts obtained by camera ISPs. Learning the forward and inverse processes of camera ISPs has been recently demonstrated, enabling physically-meaningful RAW-level image processing on input sRGB images. However, existing learning-based ISP methods fail to handle the large variations in the ISP processes with respect to camera parameters such as ISO and exposure time, and have limitations when used for various applications. In this paper, we propose ParamISP, a learning-based method for forward and inverse conversion between sRGB and RAW images, that adopts a novel neural-network module to utilize camera parameters, which is dubbed as ParamNet. Given the camera parameters provided in the EXIF data, ParamNet converts them into a feature vector to control the ISP networks. Extensive experiments demonstrate that ParamISP achieve superior RAW and sRGB reconstruction results compared to previous methods and it can be effectively used for a variety of applications such as deblurring dataset synthesis, raw deblurring, HDR reconstruction, and camera-to-camera transfer. Equal contribution. Work done prior to joining Samsung.
@@ 1 Introduction
Unlike sRGB images, RAW images provide physically-meaningful and interpretable information such as noise distributions as they have the linear relationship between image intensity and radiant energy incident on a camera sensor. Such properties of RAW images have been exploited for denoising [1, 17, 23], HDR [5, 14], and super-resolution [32, 24], leading to superior quality than using only sRGB images. However, a RAW image demands large memory to store due to the use of high-precision bits without any compression. Therefore, only sRGB images are often shared without their RAW counterparts, making it difficult to utilize the useful properties of RAW images.
Recently, several approaches have been proposed to reconstruct RAW images from sRGB images and vice versa by modeling forward and inverse ISPs [4, 30, 29, 3, 7]. However, despite notable improvements, they still suffer from limitations. First, real-world ISPs adjust their operations based on the camera parameters, e.g., exposure time and sensor sensitivity, as shown in Fig. 1. However, previous methods overlook this adaptive nature of real-world ISPs, and learn average ISP operations, which leads to low reconstruction performance. Second, previous methods adopt rather simple network architectures disregarding the complexity of the ISP operations, which leads to low RAW reconstruction quality. They construct a single network by stacking invertible or residual blocks [30, 29], or organize modules by simply arranging convolution layers [3].
@@
In this paper, we present a novel forward and inverse ISP framework, ParamISP. ParamISP is designed to faithfully reflect the real-world ISP operations that change based on camera parameters. To this end, ParamISP leverages camera parameters provided in the EXIF metadata of a JPEG image. To effectively incorporate the camera parameters, ParamISP consists of a pair of forward and inverse ISP networks that include a novel neural network module ParamNet. ParamNet extracts a feature vector from the camera parameters including exposure time, sensitivity, aperture size, and focal length, and feeds it to the forward and inverse ISP subnetworks to control their behaviors.
@@
    We propose ParamISP, a novel learning-based forward and inverse ISP framework that leverages camera parameters (e.g., exposure time, sensitivity, aperture size, and focal length) for high-quality sRGB/RAW reconstruction.
@@ 3.2 Forward ISP Network
GlobalNet globally adjusts tone and color according to the content of its input image and the optical parameter feature vector $z$. To this end, GlobalNet adopts parametric global operations and predicts their parameters. Specifically, GlobalNet adopts two types of parametric global operations: quadratic transformation $f_{q}$ and gamma correction $f_{g}$. Let us denote the input image to global adjustment operation as $I$, and its pixel as $p=(p_{r},p_{g},p_{b})$. Then, the quadratic transformation $f_{q}$ is defined as:
@@
where $f_{gq}^{n}(I)$ is defined as $f_{gq}^{n}=f_{g}^{n}(f_{q}^{n}(I))$. $f_{g}^{n}$ and $f_{q}^{n}$ are the gamma correction and quadratic transformation with the $n$-th coefficients $G^{n}$ and $W^{n}$, respectively. In our experiments, we use $N=4$ as it leads to the best quality (Refer to Tab. S3 in the supplementary material).
@@ 4.1 Ablation Study
ISP Network Architecture
@@ 5 Applications
Fig. 8 shows qualitative results using different models. Previous methods primarily focus on cyclic reconstruction. Specifically, CycleISP uses an input LDR sRGB image for the forward ISP to adjust the tone, while InvISP, composed of a single flow-based network, is not robust when intermediate RAW images are altered. On the other hand, ParamISP successfully produces an HDR reconstruction result compared to other methods. For more qualitative results, refer to the supplementary material.
Figure 8: Qualitative results of HDR reconstruction.
@@ 6 Conclusion
In this paper, we present ParamISP, a novel learning-based forward and inverse ISP framework that leverages camera parameters. To incorporate camera parameters effectively, we introduce ParamNet to control the forward and inverse ISP networks, proposing a stable and effective training strategy with respect to camera parameters. We also present novel network architectures for ISP networks that better reflect real-world ISP operations. Through our extensive experiments, ParamISP shows the state-of-the-art performance in RAW and sRGB reconstruction and the robust applicability to various applications.
Limitations and Future Work
While we have quantitatively confirmed that incorporating the aperture size and focal length improves the reconstruction quality in Tab. 2, it is still unclear how such parameters affect the behaviors of the ISP operations. While we use datasets captured by various cameras in the pre-training stage, different ISPs may behave differently with respect to different optical parameters, and such inconsistency may potentially have negative impact on a pre-trained ISP model. A more sophisticated training strategy may resolve such issue.
Acknowledgments
This work was supported by the NRF grants (RS-2023-00211658, RS-2023-00280400, 2022R1A6A1A03052954, 2023R1A2C200494611) and IITP grant (2019-0-01906, Artificial Intelligence Graduate School Program (POSTECH)) funded by the Korea government (MSIT). This work was also supported by Samsung Research Funding Center (SRFCIT1801-52) and Samsung Electronics Co.